# 11footnotetext: Corresponding authors22footnotetext: Project lead

### v1

@@ 4.2 Integral Low-rank Adaptation
To address the aforementioned challenges, we propose the integral low-rank adaption, dubbed IntLoRA, which enables PEFT in integer arithmetic. Given that backpropagation with INT types can severely impair performance, we mainly consider simulated quantization during fine-tuning, while keeping the standard INT format for storage and inference purposes. In addition, for simplicity, this work focuses solely on weight quantization. The overall pipeline of our IntLoRA is shown in Fig. 2. We first propose the Adaptation-Quantization Separation (AQS) to process the low-rank parameters, thus avoiding direct quantization on the zero-initialized $\mathbf{AB}$. Then we develop the Variance Matching Control (VMC) to adjust the distribution of the adaptation weights, followed by the uniform quantizer or log2 quantizer to obtain the quantized adaptation weights. At last, we introduce the Multiplicative Low-rank Adaptation (MLA) to adapt quantized pre-trained weights by either int-multiplication or bit-shifting, enabling independent quantizer of adaptation and pre-trained weights. More details are given below.
Adaptation-quantization separation. The vanilla LoRA adopts zero initialization on $\mathbf{AB}$ to ensure the behavior of the model is similar to the pre-trained one at the beginning of training (Hu et al., 2021). Although this initialization can improve performance, the all-zero distribution is not quantization-friendly, as validated in  Sec. 4.1 and  Sec. 5.3. To allow accurate quantization while maintaining the correct gradient, we propose the Adaptation-Quantization-Separation (AQS) strategy. The key observation is that the adaptation requires gradients from zero-initialized weights while the quantization does not. Therefore, we can split the non-zero initialized adaptation weights into the gradient-aware zero part and the gradient-free nonzero part. Formally, let $\mathbf{R}$ be the non-zero auxiliary matrix, $\mathcal{Q}$ be the quant-dequant operator, then our AQS can be formulated as:
@@
where $\mathrm{sg}(\cdot)$ denotes the stop gradient operation. Thanks to the AQS, the $\mathbf{AB}$ can be zero-initialized for the same gradient as the original LoRA, while $\mathrm{sg}({\mathbf{R}})+\mathbf{AB}$ facilitate subsequent quantization by choosing the auxiliary matrix $\mathbf{R}$. In the following part, we will ignore the $\mathrm{sg}(\cdot)$ notation for clarity.
@@
where the task-specific adaptation term is trainable and will be quantized, and the pre-trained term is already in INT and is shared across tasks. $\mathbf{I}$ is an all-one matrix. The operator $\odot$ denotes the Hadamard product of two matrices. The proposed MLA is mathematically equivalent to its additive counterpart, while is more quantization-friendly since it avoids the shared quantizer of pre-trained and adaptation weights. It is noteworthy that the adaptation term is still in FP, and we will detail its quantization strategies in Sec. 4.3.
@@
Variance matching control. One opportunity brought from the multiplicative form in Eq. 3 is that we can apply the log2-quantization on the adaptation term, thus allowing hardware-efficient bit-shifting on the pre-trained term for adaptation. However, log2-quantization is usually more difficult than common uniform quantization (Nagel et al., 2021) and requires appropriate distribution properties, e.g., most values concentrated around zero to allow for the utilization of as many quantization bins as possible on the logarithmic scale. Here, we revisit the adaptation term in Eq. 3 aiming to find useful mathematical insights. Since the $\mathbf{AB}$ is orders of magnitude smaller than $\mathbf{R}$, we analyze and approximate the adaptation term in Eq. 3 by removing $\mathbf{AB}$ from it, namely,
$s\cdot\mathbf{I}+\frac{\mathbf{R}}{\mathbf{W}_{\mathrm{round}}-z}=s\cdot\mathbf{I}+\frac{s\cdot\mathbf{R}}{s\cdot(\mathbf{W}_{\mathrm{round}}-z)}\approx s\cdot\mathbf{I}+\frac{s\cdot\mathbf{R}}{\mathbf{W-R}}=\frac{s\cdot\mathbf{W}}{\mathbf{W}-\mathbf{R}}.$ (4)
@@
From this derivation, it follows that the auxiliary matrix $\mathbf{R}$ is crucial for controlling the distribution of the adaptation term. However, there exists a dilemma in choosing an appropriate distribution for $\mathbf{R}$. On the one hand, it is desirable for the values in $\mathbf{R}$ to be larger (which are controlled by the variance of $\mathbf{R}$) as it leads to a zero-mean adaptation term, namely,
@@
$\mathbb{E}\left[\lim_{\sigma_{\mathbf{R}}\to\infty}(s\cdot\mathbf{I}+\frac{\mathbf{R}}{\mathbf{W}_{\mathrm{round}}-z})\right]=\mathbb{E}\left[\lim_{\sigma_{\mathbf{R}}\to\infty}\frac{s\cdot\mathbf{W}}{\mathbf{W}-\mathbf{R}}\right]=0,$ (5)
@@
where $\sigma_{\mathbf{R}}$ is the standard deviation of $\mathbf{R}$. On the other hand, large values in $\mathbf{R}$ can also make the term-to-be-quantized $\mathbf{W-R}$ uncorrelated to the useful $\mathbf{W}$, namely,
$\lim_{\sigma_{\mathbf{R}}\to\infty}\rho(\mathbf{W}-\mathbf{R},\mathbf{W})=\lim_{\sigma_{R}\to\infty}\frac{\sigma_{\mathbf{W}}}{\sqrt{\sigma_{\mathbf{W}}^{2}+\sigma_{\mathbf{R}}^{2}}}=0,$ (6)
@@
which results in significantly lossy reconstruction of $\mathbf{W}$ from the $\mathbf{W-R}$ using the estimator $\mathbf{W}\approx\mathcal{Q}(\mathbf{W-R})+\mathbf{R}$. Thus, values with smaller ranges in $\mathbf{R}$ and a smaller variance of $\mathbf{R}$ are expected. We give the visualization evidence of this choice dilemma of $\sigma_{\mathbf{R}}$ in Fig. 6. Due to the contradictory aspects of the desired properties of the distribution of $\mathbf{R}$, it becomes non-trivial to select a proper variance for $\mathbf{R}$. To this end, we propose the Variance Matching Control (VMC) mechanism to adjust the distribution of $\mathbf{R}$. Specifically, we first multiply $\mathbf{R}$ by the variance ratio $r=\frac{\sigma_{\mathbf{W}}}{\sigma_{\mathbf{R}}}\in\mathbb{R}^{C_{out}}$ for rough alignment from $\mathbf{R}$ to the scale of $\mathbf{W}$. After that, we further introduce a scalar $\alpha$ as an exponent of $r$, i.e., $r^{\alpha}$, to fine-grain the search for the optimal $\mathbf{R}^{*}$. As a result, the variance-matched auxiliary matrix can be denoted as $\mathbf{R}^{*}=r^{\alpha}\cdot\mathbf{R}$, and we can use this to control the distribution suitable for log2-quantization. Since $r^{\alpha}$ can be shared across tasks, it is only of negligible cost. At last, it is worth noting that in addition to the $\sigma_{\mathbf{R}}$, the shape of the distribution of $\mathbf{R}$ also has an effect on performance, and we give detailed discussion in Sec. 5.3.
@@ 4.3 Implementation of IntLoRA
Benefiting from the quantization-friendly weight distribution produced from the variance matching control, we provide two variants of our IntLoRA according to different quantizers on the adaptation term. The first variant employs the uniform quantizer on the adaptation term, thus enabling direct weight merge between the quantized adaptation term and pre-trained weights through integer multiplication. The second variant introduces log2 quantization to achieve downstream adaptation by bit-shifting the quantized pre-trained weights. More details are given below.
@@
$=\frac{1}{2^{N}}\odot\operatorname{sign}(\mathbf{V})\odot[(\mathbf{W}_{\mathrm{round}}-z)\ll(N-\mathrm{shift})]$
@@ 5.1 Experimental Setup
Baselines. Since there is little work studying PEFT of quantized diffusion models in integer arithmetic, we reproduce existing related approaches in LLMs and compare them with the proposed IntLoRA. Specifically, we include the following baselines: i) QLoRA (Dettmers et al., 2024), which fine-tunes the quantized model using FP low-rank matrix. Since the adapted weights are still in FP, we thus apply additional PTQ on the merged weights for a fair comparison. ii) QA-LoRA (Xu et al., 2023b), which uses group-shared low-rank parameters for seamless weight merge, and iii) IR-QLoRA (Qin et al., 2024), which improves the quantization with information retention as well as information elastic connection, and we also use additional PTQ on the merged FP weights.
@@
Downstream tasks. We evaluate different methods on multiple adaptation tasks, including subject-driven generation (Ruiz et al., 2023), controllable generation (Zhang et al., 2023), and style-customized image generation (Sohn et al., 2023). The subject-driven generation aims to generate images of the same subject given several images of a specific subject and a text prompt. We use the dataset released by Dreambooth (Ruiz et al., 2023) for training and testing, which contains 25 subjects with each subject corresponding to 30 prompts. For controllable generation, we consider three sub-tasks, i.e, segmentation map to image on ADE20K dataset (Zhou et al., 2017), face landmark to image on CelebA-HQ dataset (Karras, 2017), and the canny edge to image on the COCO dataset (Lin et al., 2014). For the style-customized generation, we employ the StyleDrop (Sohn et al., 2023) dataset, which includes 18 style images, and we use 6 text prompts for each style to generate images with style similar to the style image and content aligned with the text prompt.
Evaluation metrics. In the subject-driven generation, we evaluate the quality of the generated images from three crucial aspects, i.e., we use the DINO (Caron et al., 2021), CLIP-I (Radford et al., 2021) to access the subject fidelity, the CLIP-T (Radford et al., 2021) for textual prompt content fidelity, and LPIPS (Zhang et al., 2018) for sample diversity. For controllable generation task, we adopt both referenced and non-referenced image quality assessment metrics, namely FID (Heusel et al., 2017), CLIPIQA (Wang et al., 2023b), and NIQE (Mittal et al., 2012) to evaluate the performance of different methods. For style-customized generation, following StyleDrop (Sohn et al., 2023), we use CLIP-I to access the similarity between style images and generated images and use CLIP-T to evaluate the semantic fidelity between generated results and text prompts. However, as mentioned in StyleDrop (Sohn et al., 2023), these two scores are still not a perfect metric for performance evaluation, since even simply copying the style image as the generated results can also obtain a high CLIP-I score. To this end, we further adopt the ratio $\frac{\rm CLIP-I}{\rm CLIP-T}$ to comprehensively evaluate the balance between style and content fidelity.
Table 1: Quantitative comparison on subject-driven generation tasks. We mark the proposed method with a gray background. The best results are bolded.
Method 8bit-bitwidth 4bit-bitwidth
@@
LoRA-FP 0.3625 0.7431 0.2499 0.7707 0.5625 0.7431 0.2499 0.7707
QLoRA 0.3635 0.6542 0.2550 0.7959 0.1564 0.5797 0.2260 0.7824
QA-LoRA 0.3635 0.6555 0.2537 0.7952 0.3412 0.6504 0.2514 0.7950
IR-QLoRA 0.3599 0.6547 0.2542 0.7960 0.2289 0.5907 0.2307 0.7829
IntLoRAMUL 0.3929 0.6721 0.2551 0.7928 0.3479 0.6546 0.2494 0.7997
IntLoRASHIFT 0.3825 0.6641 0.2548 0.7930 0.3462 0.6478 0.2458 0.7970
Figure 3: Quantitative comparison on subject-driven generation tasks. More results are provided in Fig. 13. Zoom in for better effects.
Implementation details. We employ the StableDiffusionV1.5 (Rombach et al., 2022) as the pre-trained backbone for subject-driven generation and controllable generation. We further employ larger SDXL (Podell et al., 2023) as the pre-trained model in the style-customized generation. We use prevalent uniform affine quantization (Nagel et al., 2021) to quantize the pre-trained model weights. We append the trainable low-rank parameters on the Query, Key, Value, and Out projection in the attention layers (Vaswani et al., 2017) and keep all other layers frozen and quantized. The rank of LoRA is set to 4 for subject-driven generation and controllable generation, and 64 for style-customized generation. For compared baselines, we control the rank to ensure a fair comparison. As for the variance matching control, we employ the ratio of the maxima of the sampled distributions as a fast estimator of the variance ratio. As for the training of the quantized adaptation term, we use the Straight Through Estimator (STE) on the quantized adaptation term to allow backpropagation. Due to the page limit, we provide more details in Appendix D.
Table 2: Quantitative comparison on controllable image generation tasks.
Segmentation-to-Image Landmark-to-Face Canny-to-Image
@@
LoRA FP 31.39 0.5710 4.45 37.50 0.5645 4.80 16.05 0.5582 5.37
QLoRA 8bit 31.09 0.5669 4.44 35.73 0.5639 5.77 15.68 0.5901 5.16
QA-LoRA 8bit 31.32 0.5755 4.17 38.88 0.5667 5.56 15.34 0.6001 5.32
IR-QLoRA 8bit 31.81 0.5761 4.30 36.30 0.5683 5.74 15.70 0.6005 4.72
IntLoRAMUL 8bit 31.09 0.5746 4.47 37.98 0.5681 6.51 16.16 0.5882 5.24
IntLoRASHIFT 8bit 31.38 0.5703 4.22 34.46 0.5495 5.19 15.76 0.5913 4.83
QLoRA 4bit 71.75 0.5584 4.61 117.37 0.5578 4.84 62.49 0.5728 4.70
QA-LoRA 4bit 31.51 0.5646 4.31 43.09 0.5748 5.00 16.73 0.5745 4.52
IR-QLoRA 4bit 35.83 0.5624 4.57 39.63 0.5632 5.23 18.30 0.5733 4.58
IntLoRAMUL 4bit 30.87 0.5704 4.24 33.62 0.5667 4.55 17.58 0.5922 4.27
IntLoRASHIFT 4bit 32.85 0.5693 4.57 35.06 0.5524 5.13 17.90 0.5907 4.89
@@
Figure 4: Quantitative comparison on controllable generation tasks. More results are provided in Appendix H. Zoom in for better effects.
5.2 Main Results and Efficiency
Subject driven generation. We first use our IntLoRA to fine-tune the pre-trained models on the subject-driven generation task. This task requires a few optimization steps for fine-tuning, and the downstream weights do not deviate too much from the pre-trained capabilities. Therefore, it can reflect the short-term tuning ability of different methods. Tab. 1 shows the results under 8-bit and 4-bit weight quantization settings. It can be seen that the proposed methods consistently outperform other competitors under different bit-width stetting. For instance, the IntLoRAMUL suppresses the IR-QLoRA by 0.1190 DINO score on the 4-bit setup. Notably, the QLoRA and IR-QLoRA baseline uses additional PTQ on the fine-tuned model, leading to a significant performance drop under the 4-bit width. In contrast, even the log2-quantization of our IntLoRASHIFT works well under 4-bit thanks to the proposed VMC. We also give qualitative visualization in Fig. 3, where one can see that our IntLoRA can facilitate subject-faithful and photo-realistic image generation.
Controllable image generation. We further validate the effectiveness of IntLoRA on the controllable image generation. Since this task requires more tuning steps for adaptation in different input spaces from pre-training, it can represent the long-term adaptation ability. The results, shown in Tab. 2, indicate that our method achieves comparable results against existing strong baselines, e.g., our IntLoRAMUL outperforms the IR-QLoRA by 4.96 FID score on the 4-bit segmentation-to-image task. Fig. 4 provides the qualitative comparison, and it can be seen that the images generated by the IntLoRA-tuned model are well matched with the control signals.
Figure 5: Quantitative comparison on style-customized generation. The text prompt is “A friendly robot in [V] style”, “A panda eating bamboo in [V] style”, and “The letter ‘G’ in [V] style”, respectively. ‘Ours’ denotes the IntLoRASHIFT. More results are provided in Fig. 17.
@@ 5.2 Main Results and Efficiency
Style customized generation. Tab. 3 shows the results of our IntLoRA compared with other baselines on the style customized generation task. Taking the ratio $\rm{\frac{CLIP-I}{CLIP-T}}$ of the LoRA-FP model as the metric reference, our method achieves a similar balance ratio of style and content as the LoRA-FP. In contrast, QA-LoRA and IR-QLoRA overfit to specific style images, reflecting excessively high CLIP-I and low CLIP-T. We also provide visualization results in Fig. 5. It can be seen that our IntLoRA achieves a favorable balance between style images and text prompts, whereas some existing approaches fail, e.g., the third row and fifth column with prompt “The letter ‘G’ in [V] style”, where the generated images almost copy the original style image.
The efficiency of IntLoRA. Although the ultimate goal of all compared methods is the quantized diffusion models for fast inference, our IntLoRA can achieve this target with less cost. Specifically, as shown in Tab. 4, a clear advantage of our method lies in its seamless switch between tuning and inference, without additional PTQ. In addition, our IntLoRA is the only one that achieves INT type adaptation weights, significantly reducing the storage costs when there are many downstream tasks. Furthermore, our methods allow for weight merge using the hardware-friendly operator, naturally enabling the merged weights still in INT to facilitate the employment of edge devices.
Table 3: Quantitative comparison on style-customized image generation tasks. The $\mathrm{ratio}$ is $\frac{\mathrm{CLIP-I}}{\mathrm{CLIP-T}}$ to evaluate style-content balance.
Metrics 8bit-bitwidth 4bit-bitwidth
CLIP-I CLIP-T ratio CLIP-I CLIP-T ratio
LoRA-FP 0.6010 0.2582 2.33 0.6010 0.2582 2.33
QLoRA 0.6016 0.2583 2.33 0.5466 0.2571 2.13
QA-LoRA 0.6593 0.2381 2.77 0.6086 0.2373 2.56
IR-QLoRA 0.7025 0.2198 3.20 0.5957 0.2365 2.52
Ours-MUL 0.5832 0.2596 2.24 0.5711 0.2551 2.23
Ours-SHIFT 0.6043 0.2569 2.35 0.5708 0.2566 2.22
Table 4: Efficiency comparison under SDXL backbone 4-bit quantization setting. #param denotes the number of trainable parameters. $N$ is the downstream task number.
Methods #param(M) storage(GB) PTQ weight merge
LoRA-FP 92M 9.8+0.34N FP addtion
QLoRA 92M 1.2+0.34N FP addtion
QA-LoRA 85M 1.2+0.34N FP addition
IR-QLoRA 85M 1.2+0.34N FP addition
IntLoRAMUL 92M 1.2+0.04N int-mul
IntLoRASHIFT 92M 1.2+0.04N bit-shift

### v3

@@ 4.1 Integral Low-rank Adaptation
To address the above challenges, we propose IntLoRA to operate adaptation on the integer arithmetic. The overall pipeline is shown in Fig. 3.
Adaptation-quantization separation. The vanilla LoRA adopts zero initialization on the adaptation parameter $\mathbf{AB}$. Although this strategy can improve performance, the all-zero distribution is not quantization-friendly. To allow accurate quantization while maintaining the correct gradient, we propose the Adaptation-Quantization-Separation (AQS) mechanism. The key observation is that the adaptation requires gradients from zero-initialized weights while the quantization does not. Therefore, we can split the adaptation weights into the gradient-enabled zero part and the gradient-free nonzero part. Formally, let $\mathbf{R}$ be the auxiliary matrix to serve as the nonzero part, $\mathcal{Q}$ be the quant-dequant operator, then our AQS can be formulated as:
@@
where $\mathrm{sg}(\cdot)$ denotes the stop gradient operation. Thanks to the AQS, the $\mathbf{AB}$ can be zero-initialized for the same gradient as the original LoRA, while $\mathrm{sg}({\mathbf{R}})+\mathbf{AB}$ facilitate subsequent quantization by specifically designing the auxiliary matrix $\mathbf{R}$ as discussed in Sec. 5.4. In the following part, we will ignore the $\mathrm{sg}(\cdot)$ notation for clarity.
@@
where the task-specific adaptation term is trainable and will be quantized, and the pre-trained term is already in integer type and is shared across tasks. $\mathbf{I}$ is an all-one matrix. The operator $\odot$ denotes the Hadamard product of two matrices. The proposed MLA is mathematically equivalent to its additive counterpart, while is more quantization-friendly since it avoids the shared quantizer of pre-trained and adaptation weights. It is noteworthy that the adaptation term is still in FP16 at this step, and we will detail its quantization strategies in Sec. 4.2.
@@
Variance matching control. One opportunity brought from the multiplicative form in Eq. 3 is that we can apply the log2-quantization on the adaptation term, thus allowing more efficient bit-shifting on the pre-trained term. However, log2-quantization is notoriously more difficult than common uniform quantization (Nagel et al., 2021) and requires appropriate distribution properties, e.g., most values concentrated around zero to allow for the utilization of as many quantization bins as possible on the logarithmic scale. Here, we revisit the adaptation term in Eq. 3 aiming to find useful mathematical insights. Given the $\mathbf{AB}$ is orders of magnitude smaller than $\mathbf{R}$ (the justification is shown in Appendix F), we approximate the adaptation term in Eq. 3 by removing $\mathbf{AB}$ from it, namely,
$s\cdot\mathbf{I}+\frac{\mathbf{R}}{\mathbf{W}_{\mathrm{round}}-z}$ $=s\cdot\mathbf{I}+\frac{s\cdot\mathbf{R}}{s\cdot(\mathbf{W}_{\mathrm{round}}-z)}$ (4)
$\approx s\cdot\mathbf{I}+\frac{s\cdot\mathbf{R}}{\mathbf{W-R}}=\frac{s\cdot\mathbf{W}}{\mathbf{W}-\mathbf{R}}.$
@@
From this derivation, it follows that the auxiliary matrix $\mathbf{R}$ is crucial for controlling the distribution shape of the adaptation term. Unfortunately, we find there exists a dilemma in choosing an appropriate distribution for $\mathbf{R}$. On one hand, it is desirable for the values in $\mathbf{R}$ to be larger. Formally, let $\sigma_{\mathbf{R}}$ be the variance of the element in $\mathbf{R}$ which is a random variable, it can be derived the expectation of the adaptation term converges to zero when $\sigma_{\mathbf{R}}$ approaches infinity, namely,
@@
$\mathbb{E}\left[\lim_{\sigma_{\mathbf{R}}\to\infty}s\cdot\mathbf{I}+\frac{\mathbf{R}}{\mathbf{W-R}}\right]=\mathbb{E}\left[\lim_{\sigma_{\mathbf{R}}\to\infty}\frac{s\cdot\mathbf{W}}{\mathbf{W}-\mathbf{R}}\right]=0.$ (5)
@@
On the other hand, setting $\sigma_{\mathbf{R}}$ too large can also lead the $\mathcal{Q}(\mathbf{W-R})$ uncorrelated to the original $\mathbf{W}$, i.e, namely,
$\lim_{\sigma_{\mathbf{R}}\to\infty}\rho(\mathcal{Q}(\mathbf{W}-\mathbf{R}),\mathbf{W})=\lim_{\sigma_{R}\to\infty}\frac{\sigma_{\mathbf{W}}}{\sqrt{\sigma_{\mathbf{W}}^{2}+\sigma_{\mathbf{R}}^{2}}}=0,$ (6)
@@
where the $\rho(\cdot,\cdot)$ denotes the correlation coefficient.  Eq. 6 indicates that a over-large $\sigma_{\mathbf{R}}$ makes it difficult to reconstruct the original signal $\mathbf{W}$ through dequantizing the $\mathcal{Q}(\mathbf{W-R})$ due to the low correlation. In short, it is important to choose an appropriate $\sigma_{\mathbf{R}}$ to strike a balance between quantization difficulty and information retention. We also give the visualization of this choice dilemma in Fig. 8. To this end, we propose the Variance Matching Control (VMC) mechanism. Specifically, we first multiply $\mathbf{R}$ by the variance ratio $r=\frac{\sigma_{\mathbf{W}}}{\sigma_{\mathbf{R}}}\in\mathbb{R}^{C_{out}}$ for rough alignment from $\mathbf{R}$ to the scale of $\mathbf{W}$. After that, we introduce a scalar $\alpha$ as an exponent of $r$, i.e., $r^{\alpha}$, to fine-grain the search for the optimal $\mathbf{R}^{*}$. As a result, the variance-matched auxiliary matrix can be denoted as $\mathbf{R}^{*}=r^{\alpha}\cdot\mathbf{R}$, and we can use this to obtain the distribution suitable for log2-quantization. Since $r^{\alpha}$ can be shared across tasks, it is only of negligible cost. In addition to the $\sigma_{\mathbf{R}}$, we observe the distribution shape of $\mathbf{R}$ also has an effect on performance, and we give a detailed discussion in Sec. 5.4. It should be noted that the $\mathbf{R}$ can be online generated during fine-tuning using the distribution statistics and fixed random seed, thus avoiding the need to store its FP16 parameters.
Table 1: Quantitative comparison on subject-driven generation tasks. The notion “WxAy” represents the bit-widths of weights “W” and activations “A”. The best results are bolded.
[…]
LoRA (Hu et al., 2021) W16A16 0.4828 0.6968 0.2954 0.8076
QLoRA (Dettmers et al., 2024) W8A8 0.4153 0.6661 0.2824 0.8088
QA-LoRA (Xu et al., 2023b) W8A8 0.4156 0.6664 0.2834 0.8086
IR-QLoRA (Qin et al., 2024) W8A8 0.4070 0.6630 0.2841 0.8110
IntLoRASHIFT (Ours) W8A8 0.4353 0.6842 0.2841 0.8257
IntLoRAMUL (Ours) W8A8 0.4498 0.6882 0.2858 0.8062
QLoRA (Dettmers et al., 2024) W4A8 0.2136 0.6134 0.2510 0.8201
QA-LoRA (Xu et al., 2023b) W4A8 0.4127 0.6897 0.2700 0.8281
IR-QLoRA (Qin et al., 2024) W4A8 0.3722 0.6719 0.2707 0.8186
IntLoRASHIFT (Ours) W4A8 0.4039 0.6716 0.2709 0.8187
IntLoRAMUL (Ours) W4A8 0.4242 0.6913 0.2710 0.8181
@@ 4.2 Implementation of IntLoRA
Benefiting from the quantization-friendly weight distribution, we can implement our IntLoRA with two variants according to different quantizers on the adaptation term. The first variant employs the uniform affine quantizer on the adaptation term, thus enabling weight merge through integer-type multiplication. The second variant introduces the more hardware-friendly log2 quantizer to achieve downstream adaptation by bit-shifting the quantized pre-trained weights. More details are given below.
@@
$=\frac{1}{2^{N}}\odot\operatorname{sign}(\mathbf{V})\odot[(\mathbf{W}_{\mathrm{round}}-z)\ll(N-\mathrm{shift})],$
@@ 5.1 Experimental Setup
Datasets. We evaluate on multiple adaptation tasks, including subject-driven generation (Ruiz et al., 2023), controllable generation (Zhang et al., 2023), and style-customized image generation (Sohn et al., 2023). For the subject-driven generation, we use a subset which contains 15 text-subject pairs from the Dreambooth (Ruiz et al., 2023) dataset, for fast training and evaluation. For controllable generation, we consider three sub-tasks, i.e, Segmentation map to Image (S2I) on ADE20K dataset (Zhou et al., 2017), Landmark to Face (L2F) on CelebA-HQ dataset (Karras, 2017), and the Canny edge to Image (C2I) on the COCO dataset (Lin et al., 2014). For the style-customized generation, we employ the StyleDrop (Sohn et al., 2023) dataset, which includes 18 style images, and we use 6 text prompts for each style to generate images with style similar to the style image and content aligned with the text prompt.
Implementation details. We employ the StableDiffusionV1.5 (Rombach et al., 2022) as the pre-trained backbone for subject-driven generation and controllable generation. We further employ larger SDXL (Podell et al., 2023) as the pre-trained model in the style-customized generation. We use uniform quantization (Nagel et al., 2021) to quantize the weights per-channel and activations per-tensor. Since previous methods mainly focus on efficient training, with the tuned weights still in FP16, to make a fair comparison, we apply additional PTQ on the merged weights for efficient inference. As for the training of the quantized adaptation term, we use the Straight Through Estimator (STE) to allow back-propagation. For the proposed variance matching control, we employ the ratio of the maxima of the sampled distributions as a fast estimator for the variance. Due to the page limit, we provide more details in Appendix C.
Figure 4: Qualitative comparison on subject-driven generation tasks. More results are provided in Appendix H.
5.2 Main Results
Subject driven generation. Tab. 1 gives the results of weight-activation quantization on subject-driven generation task. It can be seen that the proposed method consistently outperforms other competitors under different bit-widths. For instance, the IntLoRAMUL suppresses the IR-QLoRA by even 0.0428 DINO score on the W8A8 setup. Notably, the QLoRA and IR-QLoRA baselines, which use additional PTQ on the merged weights, suffer a significant performance drop under the W4A8 setup. In contrast, even the challenging log2-quantization of our IntLoRASHIFT works well under W4A8. We also give qualitative visualization in Fig. 4, where one can see that our IntLoRA can facilitate subject-faithful and photo-realistic image generation.
Controllable image generation. The results of controllable image generation are shown in Tab. 2. One can see that our IntLoRA continues to outperform existing strong baselines, e.g., our IntLoRAMUL outperforms the IR-QLoRA by 4.96 FID on the 4-bit S2I setting. And it can be seen that QLoRA struggles to produce meaningful results at low 4 bit-width. We also give a qualitative comparison in Fig. 5, and it can be seen that the images generated by the IntLoRA-tuned model are well-matched with the control signals.
@@ 5.2 Main Results
methods 8-bitwidth 4-bitwidth
S2I L2F C2I S2I L2F C2I
LoRA(FP16) 31.39 37.50 16.05 31.39 37.50 16.05
QLoRA 31.09 38.88 15.34 71.75 117.37 62.49
QALoRA 31.32 38.88 15.34 31.51 43.09 16.73
IR-QLoRA 31.81 36.30 15.70 35.83 39.63 18.30
IntLoRASHIFT 31.38 34.46 15.76 32.85 35.06 17.65
IntLoRAMUL 31.08 37.52 15.26 30.87 33.62 16.32
Table 3: Comparison of training and inference efficiency with other methods. We fine-tuning the StableDiffusionV1.5 model on the Dreambooth task. The training speed is tested on one NVIDIA RTX 3090 GPU.
method nbits Training Stage Inference Stage
@@
LoRA (Hu et al., 2021) W32A32 0.68s/img 7700MB 0.6968 0.2954
QLoRA (Dettmers et al., 2024) W8A8 0.85s/img 1925MB 0.6661 0.2824
IntLoRASHIFT (Ours) W8A8 0.84s/img 1925MB 0.6842 0.2841
IntLoRAMUL (Ours) W8A8 0.87s/img 1925MB 0.6882 0.2858
QLoRA (Dettmers et al., 2024) W4A8 0.85s/img 963.1MB 0.6134 0.2510
IntLoRASHIFT (Ours) W4A8 0.84s/img 963.1MB 0.6716 0.2709
IntLoRAMUL (Ours) W4A8 0.87s/img 963.1MB 0.6913 0.2710
@@
Figure 5: Qualitative comparison on controllable generation tasks. More results are provided in Appendix H.
@@
Style customized generation. The results of the style customized generation task are shown in Fig. 6. It can be seen that our IntLoRA achieves a favorable balance between style images and text prompts, whereas some existing approaches fail. For instance, in the third row, both the QALoRA and IR-QLoRA methods directly copy the original style image under the text prompt “The letter ‘G’ in [V] style”.
5.3 Efficiency Comparison
We compare the training and inference efficiency of our IntLoRA against other baselines in Tab. 3. As for training, both IntLoRA and QLoRA only need to load quantized pre-trained weights, which is more memory efficient compared to the vanilla LoRA fine-tuning. As a result, our IntLoRA and QLoRA have similar training speeds and memory costs. However, for the inference phase, our IntLoRA can naturally obtain the quantized merged weights without additional PTQ, thus streamlining the adaptation pipeline and avoiding potential performance degradation under low bit-width. In short, compared with existing methods which only focus on training efficiency, our IntLoRA presents a both training and inference efficient paradigm.